OpenStudy (anonymous):

if \(f(x)={e^{ln~x}\over x}\), then what is \(f(1)\)?

OpenStudy (abb0t):

Are you serious?

OpenStudy (anonymous):

i'm confused because the answer should b x but thats not one of the choices..

OpenStudy (abb0t):

Why is the answer \(x\)? Don't you see that \(e^{ln(x)}\) = \(x\)

OpenStudy (abb0t):

Therefore, if you have \(\frac{x}{x}\)....

OpenStudy (anonymous):

oh wait, that would make it 1...

OpenStudy (anonymous):

omg..

OpenStudy (abb0t):

noob.

OpenStudy (anonymous):

i know.. .-.

OpenStudy (anonymous):

hey quick question..if a function is symmetric with respect to the origin that makes it symmetric to the x and y axes as well, right?

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

okay well the question says: if the curve of \(f(x)\) is symmetric with respect to the origin, then it follows that f(0)-0 f(-x)=-f(x) f(x)=f(-x) f(x) is also symmetric with respect tot he x and y-axes f(-x)=-f(-x)

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

i reposted it at http://openstudy.com/study#/updates/52144845e4b0450ed75e04fb